# Distributionally Robust Markov Games with Average Reward

### v2

@@ Appendix C Irreducible DR-MGs
Note that for $\mathsf{P}_{w}$, it holds the corresponding non-robust Bellman equation:
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (22)
We further denote that
$\mathsf{P}_{V}=\arg\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}},$ (23)
then it also holds that
$V^{\pi}_{\mathsf{P}_{V}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{V}}.$ (24)
Combining (22) and (24) implies that
$(\mathsf{P}^{\pi}_{w}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}}.$ (25)
By the definition of $\mathsf{P}_{V}$, it holds that $\mathsf{P}_{V}^{\pi}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}_{w}^{\pi}V^{\pi}_{\mathsf{P}_{w}}$, thus
$(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}},$ (26)
and it holds that
$\mathsf{P}^{\pi}_{V}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\leq(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}).$ (27)
Recursively applying it further implies that
$(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\geq(\mathsf{P}^{\pi}_{V})^{k}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}),\forall k.$ (28)
By denoting $V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}$ by $x$, the inequality above can be rewritten as
$x\geq(\mathsf{P}^{\pi}_{V})^{k}x,\forall k.$ (29)
We denote $s^{\star}\triangleq\arg\min_{s\in\mathcal{S}}x(s)$, then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s).$ (30)
By our assumption, for any $s^{\prime}\in\mathcal{S}$, there exists some $k$ such that $((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}>0$. Then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s^{\star})=x(s^{\star}),$ (31)
which implies that
$x(s^{\star})=((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+(1-((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}})x(s^{\star}).$ (32)
We then have that
$((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\star})\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime}),$ (33)
which implies $x(s^{\star})=x(s^{\prime}),\forall s^{\prime}\in\mathcal{S}$, and thus $x=me$ for some constant $m$, i.e., $V^{\pi}_{\mathsf{P}_{w}}=me+V^{\pi}_{\mathsf{P}_{V}}$.
Now we plug it in (22) and (24), and it holds that
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$ (34)
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}},$ (35)
which implies that
$g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (36)
Note that $g^{\pi}_{\mathsf{P}_{w}}\leq g^{\pi}_{\mathsf{P}_{V}}$, due to $\mathsf{P}_{w}$ is the worst-case kernel w.r.t. $g^{\pi}_{\mathsf{P}}$, and $\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}$, due to the definition of $\mathsf{P}_{V}$, thus
$0\leq g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}\leq 0.$ (37)
It further implies that
$\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$
i.e.,
$\sigma^{\pi}({V^{\pi}_{\mathsf{P}_{w}}})=\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (38)
Combine it with (22) and we have that
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\sigma^{\pi}(V^{\pi}_{\mathsf{P}_{w}}).$ (39)
Hence the Bellman equation (7) has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$, which completes the proof. ∎
@@ C.1 Convexity
$g+h(s)=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (47)
and denote any of its solutions by $(g,h)$. Additionally, consider an optimal robust policy $\pi$. Then $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\pi$:
$g+h(s)=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ (48)
@@
From Theorem 4.1, (48) is solvable, and we denote one of its solutions by $(g^{\pi},h^{\pi})$. Then we have that
$g+h(s)$ $=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (49)
$g^{\pi}+h^{\pi}(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}.$ (50)
Now from Theorem 4.1, $g=g^{\pi}=g^{\star}_{\mathcal{P}}$, thus it holds that
$h^{\pi}(s)-h(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\sum_{a}\mu(a|s)\{\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})-\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq q_{s}^{\mu}(h^{\pi}-h),$ (51)
for some policy $\mu$ and some kernel $q\in\mathcal{P}$. By the irreducibility of $q^{\mu}$, $h^{\pi}-h=ce$. Thus $(g,h)=(g^{\pi},h^{\pi}+ce)$ is also a solution to (48). ∎
Based on the equivalence between the best response policy set and the optimal policy set of the induced robust MDP, the convexity can be derived from the following result.
Lemma C.3.
Define $\Pi^{\star}\triangleq\{\pi\ |\ \text{optimal policies of }g^{\pi}_{\mathcal{P}}\}$. Then $\Pi^{\star}$ is convex.
@@
Let $\pi_{1},\pi_{2}\in\Pi^{\star}$ and define $\tilde{\pi}=\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1].$ Then by Theorem 4.1,
[…]
Alternatively, let $(g,h)$ be a solution to the optimal equation, i.e. $h(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ From Lemma C.2, this $(g,h)$ also solves the equation w.r.t. $\pi_{1},\pi_{2}$, i.e.
$g+h=r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\quad\text{and}\quad g+h=r^{\pi_{2}}+\sigma^{\pi_{2}}(h).$ (53)
It then follows that,
$g+h$ $=\alpha\big(r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\big)+(1-\alpha)\big(r^{\pi_{2}}+\sigma^{\pi_{2}}(h)\big)$ (54)
$=\sum_{a}\big(\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big)r(s,a)+\sum_{a}\big(\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big)\big(\sigma_{\mathcal{P}^{a}_{s}}(h)\big)$ (55)
$=r^{\tilde{\pi}}+\sigma^{\tilde{\pi}}(h).$ (56)
Consequently, $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\tilde{\pi}$. This implies that $g^{\tilde{\pi}}=g=g^{\star}$, and thus that $\tilde{\pi}$ is also optimal.
Hence the optimal policy set is convex, which completes the proof. ∎

### v3

@@ Appendix C Irreducible DR-MGs
Fix $\pi$ and let $P^{w}$ be any worst-case kernel for the robust average reward, so $g^{\pi}_{\mathcal{P}}=g^{\pi}_{P^{w}}$. By the irreducible average-reward MDP theory (puterman2014markov), there exists a bias $h^{\pi}_{P^{w}}$ such that
$g^{\pi}_{P^{w}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{w}{}^{\pi}(s^{\prime}\mid s)h^{\pi}_{P^{w}}(s^{\prime}),\qquad s\in\mathcal{S}.$ (27)
Now fix any vector $v\in\mathbb{R}^{\mathcal{S}}$ and define a kernel $P^{v}\in\mathcal{P}$ by taking, independently for each state $s$,
$P^{v,\pi}(\cdot\mid s)\in\arg\min_{p\in\mathcal{P}_{s}^{\pi}}p^{\top}v,$
which exists by compactness of $\mathcal{P}_{s}^{\pi}$. In particular, for $v=h^{\pi}_{P^{w}}$ we define $P^{V}\triangleq P^{h^{\pi}_{P^{w}}}$ so that
$\sigma^{\pi}_{s}(h^{\pi}_{P^{w}})\;=\;\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{w}}(s^{\prime}),\qquad s\in\mathcal{S}.$
By Assumption 1, $P^{V}{}^{\pi}$ is irreducible.
For the kernel $P^{V}$, denote by $(g^{\pi}_{P^{V}},h^{\pi}_{P^{V}})$ the corresponding non-robust gain/bias pair. It satisfies
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{V}}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{V}}(s^{\prime}),\qquad s\in\mathcal{S}.$ (28)
Rewriting (27)–(28) in vector form yields
$(\mathrm{I}-P^{w}{}^{\pi})h^{\pi}_{P^{w}}=r^{\pi}-g^{\pi}_{P^{w}}\mathbf{1},\qquad(\mathrm{I}-P^{V}{}^{\pi})h^{\pi}_{P^{V}}=r^{\pi}-g^{\pi}_{P^{V}}\mathbf{1}.$
Since $P^{w}$ is worst-case for $\pi$, we have $g^{\pi}_{P^{w}}\leq g^{\pi}_{P^{V}}$, and hence
$(\mathrm{I}-P^{w}{}^{\pi})h^{\pi}_{P^{w}}\;=\;r^{\pi}-g^{\pi}_{P^{w}}\mathbf{1}\;\geq\;r^{\pi}-g^{\pi}_{P^{V}}\mathbf{1}\;=\;(\mathrm{I}-P^{V}{}^{\pi})h^{\pi}_{P^{V}}.$ (29)
Equivalently,
$(P^{w}{}^{\pi}-\mathrm{I})h^{\pi}_{P^{w}}\;\leq\;(P^{V}{}^{\pi}-\mathrm{I})h^{\pi}_{P^{V}}.$
By the definition of $P^{V}$ as a minimizer for $h^{\pi}_{P^{w}}$ we also have, componentwise,
$P^{V}{}^{\pi}h^{\pi}_{P^{w}}\;\leq\;P^{w}{}^{\pi}h^{\pi}_{P^{w}}.$ (30)
Let $x\triangleq h^{\pi}_{P^{w}}-h^{\pi}_{P^{V}}$. Using (29) and adding/subtracting $P^{V}{}^{\pi}h^{\pi}_{P^{w}}$ we obtain
$x$ $=h^{\pi}_{P^{w}}-h^{\pi}_{P^{V}}\;\geq\;P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{V}}$
$=P^{V}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{V}}+\bigl(P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{w}}\bigr)$
$=P^{V}{}^{\pi}x+\bigl(P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{w}}\bigr).$
By (30) the last term is componentwise nonnegative, so
$x\;\geq\;P^{V}{}^{\pi}x.$
Since $P^{V}{}^{\pi}$ is irreducible and stochastic, the inequality $x\geq P^{V}{}^{\pi}x$ forces $x$ to be constant, as we prove as follows.
let $s^{\star}$ be a state with minimal coordinate $x(s^{\star})=\min_{s}x(s)$. Then for any $k\geq 1$,
$x(s^{\star})\;\geq\;\sum_{s}(P^{V}{}^{\pi})^{k}(s^{\star},s)\,x(s)\;\geq\;x(s^{\star})\sum_{s}(P^{V}{}^{\pi})^{k}(s^{\star},s)\;=\;x(s^{\star}),$
so all inequalities must be equalities. Thus whenever $(P^{V}{}^{\pi})^{k}(s^{\star},s^{\prime})>0$ we must have $x(s^{\prime})=x(s^{\star})$. By irreducibility, every state is reachable from $s^{\star}$ under some power of $P^{V}{}^{\pi}$, so $x$ is constant: $x=c\mathbf{1}$ for some $c\in\mathbb{R}$, i.e.,
$h^{\pi}_{P^{w}}=h^{\pi}_{P^{V}}+c\,\mathbf{1}.$ (31)
Using (31) in the Poisson equation (28) for $P^{V}$ gives
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s)$ $=g^{\pi}_{P^{V}}+h^{\pi}_{P^{V}}(s)+c\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{V}}(s^{\prime})+c$
$=r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{w}}(s^{\prime}).$
Equivalently,
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s).$ (32)
By construction of $P^{V}$ we have
$(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s)=\sigma^{\pi}_{s}(h^{\pi}_{P^{w}}).$
On the other hand, from the Poisson equation (27) for $P^{w}$,
$g^{\pi}_{P^{w}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+(P^{w}{}^{\pi}h^{\pi}_{P^{w}})(s)\;\geq\;r^{\pi}(s)+(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s)\;=\;g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s),$
where the inequality uses the minimizing property of $P^{V}$ for $h^{\pi}_{P^{w}}$. Thus $g^{\pi}_{P^{w}}\geq g^{\pi}_{P^{V}}$. Since $P^{w}$ is worst-case, $g^{\pi}_{P^{w}}\leq g^{\pi}_{P^{V}}$, and we conclude that $g^{\pi}_{P^{w}}=g^{\pi}_{P^{V}}$.
Combining this equality with (32), and recalling that $g^{\pi}_{P^{w}}=g^{\pi}_{\mathcal{P}}$, we obtain
$g^{\pi}_{\mathcal{P}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+\sigma^{\pi}_{s}(h^{\pi}_{P^{w}}),$
which is exactly (26) with $h=h^{\pi}_{P^{w}}$.
We then the uniqueness.
Suppose $(g,h)$ is any solution of (26). By the same construction as above, there exists some kernel $P^{V}\in\mathcal{P}$ such that
$\sigma^{\pi}_{s}(h)=(P^{V}{}^{\pi}h)(s),\quad s\in\mathcal{S}.$
Thus $(g,h)$ solves the non-robust Poisson equation for $(\pi,P^{V})$:
$g+h(s)=r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)h(s^{\prime}).$
By uniqueness of the Poisson solution (up to constants) for irreducible chains, $g=g^{\pi}_{P^{V}}$ and $h=h^{\pi}_{P^{V}}+c\mathbf{1}$ for some $c$. Since any such $P^{V}$ is also worst-case for the average reward (by the same argument as above), we have $g=g^{\pi}_{\mathcal{P}}$ and $h=h^{\pi}_{P^{w}}+c\mathbf{1}$ for some worst-case kernel $P^{w}$. This proves the claimed uniqueness up to constants. ∎
@@ C.1 Convexity
$g+h(s)=\max_{a\in\mathcal{A}}\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\},\qquad s\in\mathcal{S},$ (40)
and let $(g,h)$ be any of its solutions. Let $\pi^{\star}$ be an optimal robust policy, i.e. $\pi^{\star}\in\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$. Then $(g,h)$ also solves the robust Bellman equation associated with $\pi^{\star}$:
$g+h(s)=\sum_{a\in\mathcal{A}}\pi^{\star}(a\mid s)\bigl(r(s,a)+\sigma_{s}^{a}(h)\bigr),\qquad s\in\mathcal{S}.$ (41)
@@
By Theorem C.1, any optimal policy $\pi^{\star}$ admits a worst-case kernel $P^{\star}\in\mathcal{P}$ such that $(g^{\pi^{\star}}_{P^{\star}},h^{\pi^{\star}}_{P^{\star}})$ solves the fixed-policy equation (26) with $\pi=\pi^{\star}$. Moreover $g^{\pi^{\star}}_{P^{\star}}=g^{\star}_{\mathcal{P}}$ (the optimal robust average reward), and the optimal robust Bellman equation (40) also has a solution $(g,h)$ with $g=g^{\star}_{\mathcal{P}}$. We must show that $h$ and $h^{\pi^{\star}}_{P^{\star}}$ differ only by a constant.
Equation (40) can be written as
$g+h(s)=\max_{a}\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\}.$ (42)
Optimality of $\pi^{\star}$ implies that at each state $s$, the support of $\pi^{\star}(\cdot\mid s)$ lies in the set of maximizers of the right-hand side:
$\pi^{\star}(a\mid s)>0\;\Longrightarrow\;a\in\arg\max_{a^{\prime}}\{r(s,a^{\prime})+\sigma_{s}^{a^{\prime}}(h)\}.$
Therefore,
$g+h(s)=\sum_{a}\pi^{\star}(a\mid s)\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\},$ (43)
i.e., $(g,h)$ is a solution of the fixed-policy robust Bellman equation (41) with $\pi=\pi^{\star}$.
By Theorem C.1, the solution of the fixed-policy robust Bellman equation for $\pi^{\star}$ is unique up to constants and has gain $g=g^{\star}_{\mathcal{P}}=g^{\pi^{\star}}_{P^{\star}}$. Thus $h$ and $h^{\pi^{\star}}_{P^{\star}}$ differ only by an additive constant, and in particular $(g,h)$ is of the form
$g=g^{\star}_{\mathcal{P}},\qquad h=h^{\pi^{\star}}_{P^{\star}}+c\mathbf{1}.$
This hence completes the proof. ∎
Lemma C.3 (Convexity of optimal policies).
Define the optimal policy set
$\Pi^{\star}\triangleq\bigl\{\pi\;\big|\;g^{\pi}_{\mathcal{P}}=g^{\star}_{\mathcal{P}}\bigr\}.$
Then $\Pi^{\star}$ is convex.
@@
Let $\pi_{1},\pi_{2}\in\Pi^{\star}$ and let $\tilde{\pi}\triangleq\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1]$, i.e.
$\tilde{\pi}(a\mid s)=\alpha\pi_{1}(a\mid s)+(1-\alpha)\pi_{2}(a\mid s),\quad s\in\mathcal{S},a\in\mathcal{A}.$
Let $(g,h)$ be any solution of the optimal robust Bellman equation (40). By Lemma C.2, $(g,h)$ also solves the fixed-policy robust Bellman equations for $\pi_{1}$ and $\pi_{2}$:
$g+h(s)=r^{\pi_{1}}(s)+\sigma^{\pi_{1}}_{s}(h),\qquad g+h(s)=r^{\pi_{2}}(s)+\sigma^{\pi_{2}}_{s}(h).$
Taking the convex combination with weights $(\alpha,1-\alpha)$ yields
$g+h(s)$ $=\alpha\bigl(r^{\pi_{1}}(s)+\sigma^{\pi_{1}}_{s}(h)\bigr)+(1-\alpha)\bigl(r^{\pi_{2}}(s)+\sigma^{\pi_{2}}_{s}(h)\bigr)$
$=\sum_{a}\tilde{\pi}(a\mid s)r(s,a)+\sum_{a}\tilde{\pi}(a\mid s)\sigma_{s}^{a}(h)=r^{\tilde{\pi}}(s)+\sigma^{\tilde{\pi}}_{s}(h).$
Thus $(g,h)$ also solves the fixed-policy robust Bellman equation for $\tilde{\pi}$. By Theorem C.1, the gain of any such solution must be the robust average reward of that policy; hence
$g^{\tilde{\pi}}_{\mathcal{P}}=g.$
Since $g=g^{\star}_{\mathcal{P}}$ is the maximal robust average reward (because $\pi_{1},\pi_{2}$ are optimal), we conclude that $\tilde{\pi}$ is also optimal, i.e. $\tilde{\pi}\in\Pi^{\star}$. Therefore $\Pi^{\star}$ is convex. ∎
The convexity of the best-response sets in the multi-agent case then follows by applying Lemma C.3 to the induced single-agent DR-MDPs $M_{i}(\pi_{-i})$, as discussed.